OpenStudy (anonymous):

Use Heisenberg's Uncertainty Principle to calculate the uncertainty in the position of a 546 mg ball moving at a speed of 5.3 m/s if the speed is known to be correct within ±0.05 m/s.

OpenStudy (anonymous):

My textbook defines the Heisenberg's Uncertainty formula as: \[\Delta x \times \Delta (mv) \ge \frac{ h }{ 4\pi }\]